Need to know the LCD of a algebra equation---Y-2/5=1/6 brain has stuck a little bit.
Answer by Alan3354(69443)

Need to know the LCD of a algebra equation---Y-2/5=1/6 brain has stuck a little bit.
---------------------
Equations don't have LCD's
LCD means Least Common Denominator, so it applies to 2 or more denominators, not to equations.
----------------
If you want to solve the equation:
y-2/5=1/6
Add 2/5 to both sides
y = 1/6 + 2/5
Now you need a common denominator to add 1/6 and 2/5.
The product always works, 6*5 in this case, = 30
y = 1/6 + 2/5
y = 5/30 + 12/30
y = 17/30
That's it.